OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? 3x - 2y = 6 and 5x + 3y = -9 (4, 3) (5, 4.5) (0, -3) infinitely many solutions

OpenStudy (mertsj):

Why don't you plug these points in and see if one of them makes both sentences true?

OpenStudy (anonymous):

can u show me

OpenStudy (mertsj):

\[3x-2y = 6 and 5x+3y=-9\]

OpenStudy (anonymous):

(0,-3)

OpenStudy (mertsj):

Oh well, now Kija wants to help you. Perhaps she will show you how to substitute the points.

OpenStudy (anonymous):

yes I will

OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? x + y = -8 and x - y = 8 (0, -8) (-12, 4) (-4, -4) no solution

OpenStudy (anonymous):

with this can u show me step by step how to do it

OpenStudy (anonymous):

Yea sure

OpenStudy (anonymous):

Okay so lets start with the first choice: (0,-8) Substitute it in: 0 -8 = -8 and 0 +8 = 8 Does it work?

OpenStudy (anonymous):

it cancels everything out right

OpenStudy (anonymous):

Right so we have our answer right there we don't even have to figure out the rest of them.

OpenStudy (anonymous):

oh i see thank you i understand

OpenStudy (anonymous):

np :)